From: Frank Geurts (geurts_at_rice.edu)
Date: Mon Mar 17 2003 - 10:52:01 EST
yep, tof ran only for a brief period between 23h11 and 2h18 this morning
... like before no apparent problems from the daq-operation.
More later.
-frank
Tonko A. Ljubicic wrote:
> On Mon, 17 Mar 2003, W.J. Llope wrote:
>
>
>>...ran fine. data looks A.O.K.....
>>
>>looks like eleanor's reseating of a lemo cable at the TOF TCD
>>did the trick... thanks again eleanor..
>
>
> Hm, according to the shift-leader, it did NOT run fine and as such is
> NOT in the run... I don't know anything more...
>
> -- Tonko
>
>
> --
> Tonko (Ante) Ljubicic Email: tonko_at_bnl.gov
> Bldg. 510A Tel: +1-631-344-7346
> Brookhaven National Lab., P.O. Box 5000 Fax: +1-631-344-4206
> Upton, NY, 11973
>
This archive was generated by hypermail 2.1.4 : Thu Jul 24 2003 - 00:39:38 EDT